OpenStudy (anonymous):

can someone please explain to me how to solve this problem? 5(7-x) + 6x(3x+4)

OpenStudy (anonymous):

I meant how to simplify it:)

OpenStudy (anonymous):

yea you want to distribute the outside numbers so 5(7-x) = 35 - 5x the 6x(3x+4) = 18x^2 +24x so combine that and you get: 18x^2+ 19x + 35

OpenStudy (anonymous):

can you simplify any more than that?

OpenStudy (anonymous):

that's fully simplified

OpenStudy (anonymous):

ok thanx I just thought that the 18x^2 and the 19x could be combined but I'm probably just remembering it wrong:)

OpenStudy (coolshubhs):

solve it using long division method..